# Enhanced Feature Learning via Regularisation: Integrating Neural Networks and Kernel Methods

### v1

@@ 4.2 Bound on Expected Risk of Regularised Estimator
We now use the bounds on the Gaussian complexity we have obtained in Section 4.1 to derive a bound on the expected risk of BKerNN. We show that, with explicit rates, the expected risk of our estimator converges with high-probability to that of the minimiser for data with subgaussian norms, which includes both bounded data and data with subgaussian components.
First, we provide a definition of subgaussian real variables, as given by Vershynin (2018).
@@
    Well-specified model: The minimiser $f^{*}:=\operatorname*{arg\,min}_{f\in\mathcal{F}_{\infty},\Omega(f)<+\infty}\mathcal{R}(f)$ exists.
@@
    Convexity of the loss: For any $(x,y)\in\mathcal{X}\times\mathcal{Y}$. $f\in\mathcal{F}_{\infty}\to\ell(y,f(x))$ is convex.
@@
$G_{n}\leq\min\bigg{(}\frac{3}{n^{1/6}}\big{(}(\log 2d)^{1/4}\mathds{1}_{*=\infty}+\mathds{1}_{*=2}\big{)}\bigg{(}\mathbb{E}_{\mathcal{D}_{n}}\big{(}\max_{i\in[n]}(\|X_{i}\|^{*})^{2}\big{)}\bigg{)}^{1/4}$ ,
@@
Proof [Theorem 16] This proof is primarily based on Bach (2024, Proposition 4.7).
@@
If the model is not well-specified but we consider the Bayes predictor $f^{*}$ to be Lipschitz continuous, the rates for neural networks with ReLU activation and bounded corresponding Banach norm $\gamma_{1}$ ($O(n^{-1/(d+5)})$) and kernel methods ($O(n^{-1/(d+1)})$) (Bach, 2024, Section 7.5, Section 9.4) do not beat the curse of dimensionality, and neither does our setup.
However, in the case of linear latent variables, i.e., under the multiple index model where $f^{*}=g^{*}(P^{\top}x)$ with $P$ a $d\times k$ matrix with $k<d$ and orthonormal columns, the RKHS cannot take advantage of this hypothesis and the rates remain unchanged. In contrast, the neural network can, assuming that $g^{*}$ has bounded Banach norm, then we only pay the price of the $k$ underlying dimensions and not the full $d$ dimensions (Bach, 2024, Section 9.4). BKerNN also has this property, which is visible by using the simple arguments presented in the discussion in Bach (2024, Section 9.3.5), which show that $\Omega(f)\leq\Omega(g^{*})$. Moreover, the optimisation process for BKerNN is much easier than that of neural networks, and our function space is larger, underscoring the attractiveness of BKerNN.
@@ 5.1 Introduction to Scores and Competitors
In the experiments below, we use two scores to assess performance. The prediction score is defined as the coefficient of determination, a classical metric in the statistics literature (Wright, 1921), $R^{2}$, which ranges from $\infty$- ∞ to 1, where a score of 1 indicates perfect prediction, a score of 0 indicates that the model predicts no better than the mean of the target values, and negative values indicate that the model performs worse than this baseline. Mathematically, the $R^{2}$ score is defined as follows:
@@ 5.2 Experiment 1: Optimisation Procedure, Importance of Positive Homogeneous Kernel
Figure 1: MSE across optimisation procedure for different kernels
@@ 5.3.2 Experiment 3
Figure 2: Influence of parameters: left: $m$, middle: $\lambda$, right: type of penalty
@@
Figure 3: Comparison to neural network on 1D examples
@@
Figure 4: Performance comparison across varying sample sizes and dimensions
@@
In Experiment 6, we evaluate the $R^{2}$ scores, defined in Equation (15), of four methods: BKRR, BKerNN with concave variable regularisation, BKerNN with concave feature regularisation, and ReLUNN, across 17 real-world data sets. These data sets were obtained from the tabular benchmark numerical regression suite via the OpenML platform, as described by Grinsztajn et al. (2022). Each data set was processed to include only numerical variables and rescaled to have centred covariates with standard deviation equal to one. The data sets were uniformly cropped to contain 400 training samples and 100 testing samples, with dimensionality varying across data sets as shown in Figure 5. For both BKerNN and ReLUNN, the number of particles or hidden neurons was set to twice the dimension of each data set, while the training parameters were fixed. Details are available in Appendix B.5.
@@
Figure 5: Comparison of $R^{2}$ scores on real data sets
@@ 6 Conclusion
To conclude, we have introduced a novel framework for feature learning and function estimation in supervised learning, termed Brownian kernel neural network (BKerNN). By leveraging regularised empirical risk minimisation over averages of Sobolev spaces on one-dimensional projections of the data, we established connections to kernel ridge regression and infinite-width one-hidden layer neural networks. We provide an efficient computational method for BKerNN, emphasising the importance of the positive homogeneity of the Brownian kernel. Through rigorous theoretical analysis, we demonstrated that, in the well-specified setting for subgaussian data, BKerNN achieves convergence of its expected risk to the minimal risk with explicit rates, potentially independent of the data dimension, underscoring the efficacy of our approach. We have extensively discussed the relationship between the space of functions we propose and other classical functions spaces. Numerical experiments across simulated scenarios and real data sets confirm BKerNN’s superiority over traditional kernel ridge regression and competitive performance with neural networks employing ReLU activations, achieved with fewer particles or hidden neurons. Future research directions include the development of more efficient algorithms for the computation of the estimator, improved analysis of the Gaussian complexity, and theoretical investigation of other penalties.
@@ A.1.1 Proof of Lemma 3
Proof [Lemma 3] We first check that $\mathcal{F}_{\infty}$ is a vector space.
Let $f\in\mathcal{F}_{\infty}$ with $f(\cdot)=c+\int_{\mathcal{S}^{d-1}}g_{w}(w^{\top}\cdot){\rm d}\mu(w)$ and $\tau\in\mathbb{R}$ then $\tau f(\cdot)=\tau c+\int_{\mathcal{S}^{d-1}}\tau g_{w}(w^{\top}\cdot){\rm d}\mu(w)$ and $\tau g_{w}\in\mathcal{H}$ hence $\tau f\in\mathcal{F}_{\infty}$.
Now let $f,\tilde{f}\in\mathcal{F}_{\infty}$, then $(f+\tilde{f})(\cdot)=c+\tilde{c}+\int_{\mathcal{S}^{d-1}}g_{w}(w^{\top}\cdot){\rm d}\mu(w)+\int_{\mathcal{S}^{d-1}}\tilde{g}_{w}(w^{\top}\cdot){\rm d}\tilde{\mu}(w)=c+\tilde{c}+\int_{\mathcal{S}^{d-1}}({\rm d}\frac{2\mu(w)}{\mu(w)+\tilde{\mu}(w)}g_{w}+{\rm d}\frac{2\tilde{\mu}(w)}{\mu(w)+\tilde{\mu}(w)}\tilde{g}_{w})(w^{\top}\cdot){\rm d}(\frac{\mu+\tilde{\mu}}{2})(w)$, hence $f+\tilde{f}$ belongs to $\mathcal{F}_{\infty}$ and $\mathcal{F}_{\infty}$ is a vector space.
@@
Then, we see as a direct consequence that $\{f\in\mathcal{F}_{\infty},\Omega_{0}(f)<\infty\}=\{f\in\mathcal{F}_{\infty},\Omega(f)<\infty\}$ is also a vector space. Let $f,\tilde{f}\in\mathcal{F}_{\infty}$ with $\Omega_{0}(f)<\infty$ and $\tau\in\mathbb{R}$. Then, $\Omega(f)=0\iff f=0$ and $\Omega(\tau f)=|\tau|\Omega(f)$. For the triangular inequality, we have
$\Omega_{0}(f+\tilde{f})$ $=\int_{\mathcal{S}^{d-1}}\bigg{\|}{\rm d}\frac{2\mu(w)}{\mu(w)+\tilde{\mu}(w)}g_{w}+{\rm d}\frac{2\tilde{\mu}(w)}{\mu(w)+\tilde{\mu}(w)}\tilde{g}_{w}\bigg{\|}_{\mathcal{H}}{\rm d}\bigg{(}\frac{\mu+\tilde{\mu}}{2}\bigg{)}(w)$
$\leq\int_{\mathcal{S}^{d-1}}{\rm d}\frac{2\mu(w)}{\mu(w)+\tilde{\mu}(w)}\|g_{w}\|_{\mathcal{H}}+{\rm d}\frac{2\tilde{\mu}(w)}{\mu(w)+\tilde{\mu}(w)}\|\tilde{g}_{w}\|_{\mathcal{H}}{\rm d}\bigg{(}\frac{\mu+\tilde{\mu}}{2}\bigg{)}(w)$
$\leq\Omega_{0}(f)+\Omega_{0}(\tilde{f}),$
hence $\Omega(f+\tilde{f})=\max(f(0)+\tilde{f}(0),\Omega_{0}(f+\tilde{f}))\leq\Omega(f)+\Omega(\tilde{f})$, which yields that $\Omega$ is a norm.
We first check the Hölder continuous property
$f(x)-f(x^{\prime})$ $=c+\int_{\mathcal{S}^{d-1}}g_{w}(w^{\top}x){\rm d}\mu(w)-c-\int_{\mathcal{S}^{d-1}}g_{w}(w^{\top}x^{\prime}){\rm d}\mu(w)$
$=\int_{\mathcal{S}^{d-1}}\langle g_{w},k_{w^{\top}x}-k_{w^{\top}x^{\prime}}\rangle{\rm d}\mu(w)$
$|f(x)-f(x^{\prime})|$ $\leq\int_{\mathcal{S}^{d-1}}\|g_{w}\|_{\mathcal{H}}\|k_{w^{\top}x}-k_{w^{\top}x^{\prime}}\|_{\mathcal{H}}{\rm d}\mu(w)\leq\int_{\mathcal{S}^{d-1}}\|g_{w}\|_{\mathcal{H}}\sqrt{|w^{\top}(x-x^{\prime})|}{\rm d}\mu(w)$
$\leq\int_{\mathcal{S}^{d-1}}\|g_{w}\|_{\mathcal{H}}\sqrt{\|x-x^{\prime}\|^{*}}{\rm d}\mu(w)\leq\Omega_{0}(f)\sqrt{\|x-x^{\prime}\|^{*}}.$
We then consider the weak derivatives, let $a\in[d]$, then
$\frac{\partial f}{\partial x_{a}}(x)=\int_{\mathcal{S}^{d-1}}w_{a}g_{w}^{\prime}(w^{\top}x){\rm d}\mu(w),$
which means that
$\int_{\mathbb{R}^{d}}\left(\frac{\partial f}{\partial x_{a}}(x)\right)^{2}{\rm d}\rho(x)$ $=\int_{\mathbb{R}^{d}}\left(\int_{\mathcal{S}^{d-1}}w_{a}g_{w}^{\prime}(w^{\top}x){\rm d}\mu(w)\right)^{2}{\rm d}\rho(x)$
$\leq\int_{\mathbb{R}^{d}}\left(\int_{\mathcal{S}^{d-1}}|w_{a}||g_{w}^{\prime}(w^{\top}x)|{\rm d}\mu(w)\right)^{2}{\rm d}\rho(x)$
$\leq\int_{\mathbb{R}^{d}}\left(\int_{\mathcal{S}^{d-1}}w_{a}^{2}g_{w}^{\prime}(w^{\top}x)^{2}{\rm d}\mu(w)\int_{\mathcal{S}^{d-1}}{\rm d}\mu(w)\right){\rm d}\rho(x)$
$=\int_{\mathcal{S}^{d-1}}\int_{\mathbb{R}^{d}}w_{a}^{2}g_{w}^{\prime}(w^{\top}x)^{2}{\rm d}\rho(x){\rm d}\mu(w)$
$=\int_{\mathcal{S}^{d-1}}w_{a}^{2}\int_{\mathbb{R}^{d}}g_{w}^{\prime}(w^{\top}x)^{2}{\rm d}\rho(x){\rm d}\mu(w)$
$\leq\int_{\mathcal{S}^{d-1}}w_{a}^{2}\|g_{w}\|_{\mathcal{H}}{\rm d}\mu(w)\leq\Omega_{0}(f)\sup_{w\in\mathcal{S}^{d-1}}w_{a}^{2}$
[…]

### v2

@@ 4.2 Bound on Expected Risk of Regularised Estimator
We now use the bounds on the Gaussian complexity we have obtained in Section 4.1 to derive a bound on the expected risk of BKerNN. We show that, with explicit rates, the expected risk of our estimator converges with high-probability to that of the minimiser for data with subgaussian norms, which includes both bounded data and data with subgaussian components. First, we provide a definition of subgaussian real variables, as given by Vershynin (2018).
@@
    Well-specified model: The minimiser $f^{*}:=\operatorname*{arg\,min}_{f\in\mathcal{F}_{\infty}}\mathcal{R}(f)$ exists.
@@
    Convexity of the loss: For any $(x,y)\in\mathcal{X}\times\mathcal{Y}$, $f\in\mathcal{F}_{\infty}\to\ell(y,f(x))$ is convex.
@@
$G_{n}\leq\min\bigg{(}\frac{{\color[rgb]{0,0,0}6}}{n^{1/6}}\big{(}(\log 2d)^{1/4}\mathds{1}_{*=\infty}+\mathds{1}_{*=2}\big{)}\bigg{(}\mathbb{E}_{\mathcal{D}_{n}}\big{(}\max_{i\in[n]}(\|X_{i}\|^{*})^{2}\big{)}\bigg{)}^{1/4}$ ,
@@
Proof [Proof of Theorem 16] This proof is primarily based on Bach (2024, Proposition 4.7).
@@
However, in the case of linear latent variables, i.e., under the multiple index model where $f^{*}=g^{*}(P^{\top}x)$ with $P$ a $d\times k$ matrix with $k<d$ and orthonormal columns, the RKHS cannot take advantage of this hypothesis and the rates remain unchanged. In contrast, the neural network can, assuming that $g^{*}$ has bounded Banach norm, then we only pay the price of the $k$ underlying dimensions and not the full $d$ dimensions (Bach, 2024, Section 9.4). BKerNN also has this property, which is visible by using the simple arguments presented in the discussion in Bach (2024, Section 9.3.5), which show that $\Omega(f^{*})\leq\Omega(g^{*})$. Moreover, the optimisation process for BKerNN is much easier than that of neural networks, and our function space is larger, underscoring the attractiveness of BKerNN. Our analysis in this section could be complemented by an analysis of the mis-specified case (i.e., $f^{\ast}$ only assumed to be Lipschitz), where, controlling both approximation and estimation errors, we should expect our method to benefit from the same adaptivity to linear latent variables as neural networks (Bach, 2017).
@@ 5.1 Introduction to Scores and Competitors
In the experiments below, we use two scores to assess performance. The prediction score is defined as the coefficient of determination, a classical metric in the statistics literature (Wright, 1921), $R^{2}$, which ranges from $\infty$- ∞ to 1, where a score of 1 indicates perfect prediction, a score of 0 indicates that the model predicts no better than the mean of the target values, and negative values indicate that the model performs worse than this baseline. Mathematically, the $R^{2}$ score is defined as follows
@@ 5.2 Experiment 1: Optimisation Procedure, Importance of Positive Homogeneous Kernel
Figure 1: MSE across optimisation procedure for different kernels.
@@ 5.3.2 Experiment 3
Figure 2: Influence of parameters: left: $m$, middle: $\lambda$, right: type of penalty.
@@
Figure 3: Comparison to neural network on 1D examples.
@@
Figure 4: Performance comparison across varying sample sizes and dimensions.
@@
In Experiment 6, we evaluate the $R^{2}$ scores, defined in Equation (15), of four methods: BKRR, BKerNN with concave variable regularisation, BKerNN with concave feature regularisation, and ReLUNN, across 17 real-world data sets. These data sets were obtained from the tabular benchmark numerical regression suite via the OpenML platform, as described by Grinsztajn et al. (2022). Each data set was processed to include only numerical variables and rescaled to have centred covariates with standard deviation equal to one. The data sets were uniformly cropped to contain 400 training samples and 100 testing samples (except for a few datasets, see Appendix B.5), with dimensionality varying across data sets as shown in Figure 5. For both BKerNN and ReLUNN, the number of particles or hidden neurons was set to twice the dimension of each data set (except for the dataset with the highest dimension, “semeion”, $d=256$, where it was fixed to a 100), while the training parameters were fixed. Details are available in Appendix B.5.
@@
Figure 5: Comparison of $R^{2}$ scores on real data sets.
@@ 6 Conclusion
To conclude, we have introduced a novel framework for feature learning and function estimation in supervised learning, termed Brownian kernel neural network (BKerNN). By leveraging regularised empirical risk minimisation over averages of Sobolev spaces on one-dimensional projections of the data, we established connections to kernel ridge regression and infinite-width one-hidden layer neural networks. We provide a particle-based computational method for BKerNN, emphasising the importance of the positive homogeneity of the Brownian kernel. Through rigorous theoretical analysis, we demonstrated that, in the well-specified setting for subgaussian data, BKerNN achieves convergence of its expected risk to the minimal risk with explicit rates, potentially independent of the data dimension, underscoring the efficacy of our approach. We have extensively discussed the relationship between the space of functions we propose and other classical functions spaces. Numerical experiments across simulated scenarios and real data sets confirm BKerNN’s superiority over traditional kernel ridge regression and competitive performance with neural networks employing ReLU activations, achieved with fewer particles or hidden neurons. Future research directions include the development of more efficient algorithms for the computation of the estimator, improved analysis of the Gaussian complexity, and theoretical investigation of other penalties.
@@ A.1.1 Proof of Lemma 3
Proof [Proof of Lemma 3] We first check that $\mathcal{F}_{\infty}$ is a vector space.
Let $f\in\mathcal{F}_{\infty}$, a constant $c$ and a sequence of measures $(\mu_{i})_{i\in\mathbb{N}}$ defining $f$ (i.e. $\forall i\in\mathbb{N},f=c+\int_{\mathcal{H}\times\mathcal{S}^{d-1}}g(w^{\top}\cdot){\rm d}\mu_{i}(g,w))$ such that $\Omega_{0}(f)=\lim_{i\to\infty}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|g\|_{\mathcal{H}}{\rm d}\mu_{i}(g,w)$. Recall that $c$ is unique for a given $f$ as it is equal to $f(0)$.
Then, let $\tau\in\mathbb{R}$. We have $\forall i\in\mathbb{N},\tau f=\tau c+\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\tau g(w^{\top}\cdot){\rm d}\mu_{i}(g,w)=\tau c+\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\tilde{g}(w^{\top}\cdot){\rm d}\tilde{\mu_{i}}(\tilde{g},w)$, with the change of variable $\tilde{g}=\tau g$ inducing the change on the measure ${\rm d}\tilde{\mu_{i}}(\tilde{g},w)={\rm d}\mu_{i}(\tilde{g}/\tau,w)$ and $\tilde{\mu_{i}}$ is still a probability measure (using the pushforward measure technique, see Peyré and Cuturi (2019)) . We also see that for any $i$,
×$\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|\tilde{g}\|_{\mathcal{H}}{\rm d}\tilde{\mu_{i}}(\tilde{g},w)=|\tau|\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|g\|_{\mathcal{H}}{\rm d}\mu_{i}(g,w),$
and hence since $|\tau|\Omega_{0}(f)=|\tau|\lim_{i\to\infty}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|g\|_{\mathcal{H}}{\rm d}\mu_{i}(g,w)$, we have
$\lim_{i\to\infty}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|\tilde{g}\|_{\mathcal{H}}{\rm d}\tilde{\mu_{i}}(\tilde{g},w)=|\tau|\Omega_{0}(f)$
and therefore $\Omega_{0}(\tau f)<\infty$ and $\Omega_{0}(\tau f)\leq|\tau|\Omega_{0}(f)$, which means that $\tau f$ belongs to $\mathcal{F}_{\infty}$. Now we can also use this property to obtain that $\Omega_{0}(f)=\Omega_{0}(\frac{1}{\tau}\tau f)\leq\frac{1}{|\tau|}\Omega_{0}(\tau f)\leq\Omega_{0}(f)$ and hence $\Omega_{0}(\tau f)$ is actually equal to $|\tau|\Omega_{0}(f)$.
Now let $f_{1},f_{2}\in\mathcal{F}_{\infty}$, defined respectively using $c_{1},c_{2}\in\mathbb{R}$ and two sequences of measures $(\mu_{i}^{(1)})_{i\in\mathbb{N}},(\mu_{i}^{(2)})_{i\in\mathbb{N}}$ such that $\Omega_{0}(f_{1})=\lim_{i\to\infty}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|g\|_{\mathcal{H}}{\rm d}\mu_{i}^{(1)}(g,w)$ and $\Omega_{0}(f_{2})=\lim_{i\to\infty}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|g\|_{\mathcal{H}}{\rm d}\mu_{i}^{(2)}(g,w)$. Now for any $i\in\mathbb{N}$, we have
$(f_{1}+f_{2})(\cdot)=c_{1}+c_{2}+\int_{\mathcal{H}\times\mathcal{S}^{d-1}}g(w^{\top}\cdot){\rm d}\mu_{i}^{(1)}(g,w)+\int_{\mathcal{H}\times\mathcal{S}^{d-1}}g(w^{\top}\cdot){\rm d}\mu_{i}^{(2)}(g,w),$
and hence $(f_{1}+f_{2})(\cdot)=(c_{1}+c_{2})+\int_{\mathcal{H}\times\mathcal{S}^{d-1}}g(w^{\top}\cdot){\rm d}\left(\mu_{i}^{(1)}+\mu_{i}^{(2)}\right)(g,w)$ and moreover, $(f_{1}+f_{2})(\cdot)=(c_{1}+c_{2})+\int_{\mathcal{H}\times\mathcal{S}^{d-1}}2g(w^{\top}\cdot){\rm d}\left(\frac{\mu_{i}^{(1)}+\mu_{i}^{(2)}}{2}\right)(g,w)$.
Using a change of variable as we did for $\tau f$, this means that we can write $f_{1}+f_{2}$ using a probability measure and a real constant, in the same way as the functions in $\mathcal{F}_{\infty}$. For $f_{1}+f_{2}$ to belong to $\mathcal{F}_{\infty}$, we now need to check that $\Omega_{0}(f_{1}+f_{2})$ is well defined.
We first consider the function $\tilde{f}:=\int_{\mathcal{H}\times\mathcal{S}^{d-1}}g(w^{\top}\cdot){\rm d}\left(\frac{\mu_{i}^{(1)}+\mu_{i}^{(2)}}{2}\right)(g,w)$ (for any $i\in\mathbb{N})$. We have that
$\lim_{i\to\infty}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|g\|_{\mathcal{H}}{\rm d}\left(\frac{\mu_{i}^{(1)}+\mu_{i}^{(2)}}{2}\right)(g,w)=\frac{\Omega_{0}(f_{1})+\Omega_{0}(f_{2})}{2},$
by simply splitting the integral. Therefore the function $\tilde{f}$ is such that $\Omega_{0}(\tilde{f})<\infty$ and belongs to $\mathcal{F}_{\infty}$ with $\Omega_{0}(\tilde{f})\leq\frac{\Omega_{0}(f_{1})+\Omega_{0}(f_{2})}{2}$. It suffices then to notice that $f_{1}+f_{2}=c_{1}+c_{2}+2\tilde{f}$ (and that the intercept has no impact on the definition of $\Omega_{0}$) to obtain that $\Omega_{0}(f_{1}+f_{2})$ is well defined and that $\Omega_{0}(f_{1}+f_{2})\leq\Omega_{0}(f_{1})+\Omega_{0}(f_{2})$, hence $f_{1}+f_{2}$ belongs to $\mathcal{F}_{\infty}$.
@@
This yields that $\mathcal{F}_{\infty}$ is a vector space. We also see that $\Omega(f_{1}+f_{2})=\max(f_{1}(0)+f_{2}(0),\Omega_{0}(f_{1}+f_{2}))\leq\Omega(f_{1})+\Omega(f_{2})$. Since $\Omega(f)=0\iff f=0$, we have that $\Omega$ is a norm on $\mathcal{F}_{\infty}$.
We now check the Hölder continuity property. Let $\mu$ be any measure defining $f\in\mathcal{F}_{\infty}$ and $c=f(0)$, then
$f(x)-f(x^{\prime})$ $=c+\int_{{\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}}{\color[rgb]{0,0,0}g}(w^{\top}x){\rm d}\mu({\color[rgb]{0,0,0}g,w})-c-\int_{{\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}}{\color[rgb]{0,0,0}g}(w^{\top}x^{\prime}){\rm d}\mu({\color[rgb]{0,0,0}g,w})$
=×$=\int_{{\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}}\langle{\color[rgb]{0,0,0}g},k_{w^{\top}x}-k_{w^{\top}x^{\prime}}\rangle{\rm d}\mu({\color[rgb]{0,0,0}g,w})$
$|f(x)-f(x^{\prime})|$$\leq\int_{{\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}}\|{\color[rgb]{0,0,0}g}\|_{\mathcal{H}}\|k_{w^{\top}x}-k_{w^{\top}x^{\prime}}\|_{\mathcal{H}}{\rm d}\mu({\color[rgb]{0,0,0}g,w})$
×$\leq\int_{{\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}}\|{\color[rgb]{0,0,0}g}\|_{\mathcal{H}}\sqrt{|w^{\top}(x-x^{\prime})|}{\rm d}\mu({\color[rgb]{0,0,0}g,w})$
×$\leq\int_{{\color[rgb]{0,0,0}\mathcal{H}\times\mathcal{S}^{d-1}}}\|{\color[rgb]{0,0,0}g}\|_{\mathcal{H}}\sqrt{\|x-x^{\prime}\|^{*}}{\rm d}\mu({\color[rgb]{0,0,0}g,w})\leq\Omega_{0}(f)\sqrt{\|x-x^{\prime}\|^{*}}.$